# BabyVLM: Data-Efficient Pretraining of VLMs Inspired by Infant Learningthanks: Project website: shawnking98.github.io/BabyVLM

### v1

@@ Abstract
Human infants rapidly develop visual reasoning skills from minimal input, suggesting that developmentally inspired pretraining could significantly enhance the efficiency of vision-language models (VLMs). Although recent efforts have leveraged infant-inspired datasets like SAYCam, existing evaluation benchmarks remain misaligned—they are either too simplistic, narrowly scoped, or tailored for large-scale pretrained models. Additionally, training exclusively on infant data overlooks the broader, diverse input from which infants naturally learn. To address these limitations, we propose BabyVLM, a novel framework comprising comprehensive in-domain evaluation benchmarks and a synthetic training dataset created via child-directed transformations of existing datasets. We demonstrate that VLMs trained with our synthetic dataset achieve superior performance on BabyVLM tasks compared to models trained solely on SAYCam or general-purpose data of the SAYCam size. BabyVLM thus provides a robust, developmentally aligned evaluation tool and illustrates how compact models trained on carefully curated data can generalize effectively, opening pathways toward data-efficient vision-language learning paradigms.
@@ 1 Introduction
We propose a novel framework, BabyVLM, for data-efficient pretraining of vision-language models (VLMs). To this end we introduce methods for creating minimal yet naturalistic data—akin to the input human infants receive—as well as comprehensive in-domain evaluation benchmarks. By carefully curating the training data, we show that our method yields more robust, baby-like representations compared to training on general-purpose corpora, and can further serves as a template for resource-efficient model training in other specialized domains.
Challenges in Current VLM Training. Vision-Language Models have advanced rapidly in recent years [35, 46, 15, 55, 22], but these advancements often rely on massive datasets and prohibitively expensive computational resources. For instance, training large-scale models such as LLaMA or LLaVA can require thousands of GPU hours [49, 14, 43, 22]. Such demands pose fundamental barriers for independent researchers with limited resources, highlighting the need for more accessible pretraining methods.
@@
  • In-Domain Evaluation Tasks. We design three novel evaluation tasks derived from the SAYCam dataset. These tasks are tailored to reflect the cognitive and perceptual abilities typical of early human development, enabling comprehensive and meaningful evaluation of compact models trained on developmentally plausible data.
@@
Figure 2: Pipeline for generating the transferred dataset. Step 1: We prompt GPT-4o to check whether an input caption is describing something a child would see in daily life, and transfer the original image captions into simpler, child-directed utterances. Step 2: We use CLIP similarity score as a metric to represent the distance between two images, then conduct Hungarian matching to select a small subset of the transferred dataset that is visually aligned with SAYCam images.
@@ 2 Related Work
Vision-Language Models. Large vision-language models (VLMs) [35, 22, 15, 46, 4, 25] have significantly advanced multimodal understanding by integrating visual and linguistic data for various tasks, including image captioning, visual question answering, and conversational interaction. Early influential models such as CLIP [35] leveraged contrastive learning paradigms, effectively aligning visual and textual representations within a unified embedding space. More recent generative frameworks, such as LLaVA [22], have combined pretrained visual encoders [35] with large language models [60], enabling more advanced conversational interactions and multimodal generative capabilities. However, these models typically require extensive computational resources and large-scale datasets. In contrast, our approach specifically targets compact generative VLMs trained exclusively on developmentally plausible datasets, providing a framework to improve data efficiency and better align model training with cognitive development processes observed in human infants.
@@
Multimodal Benchmarks. Existing multimodal evaluation benchmarks can be broadly classified as general-purpose or developmentally inspired. General-purpose benchmarks, such as Visual Question Answering (VQA) [2, 57, 3] and Winoground [48], evaluate advanced visio-linguistic integration but typically rely on large-scale, non-developmental datasets, rendering them unsuitable for compact models trained on limited developmental data. Conversely, developmentally inspired benchmarks—such as Labeled-S [31], ModelVsBaby [41], DevBench [45], and MEWL [16]—are more aligned with early cognitive processes but often limited to simplistic classification tasks or utilize data not fully reflective of the training domain. Our work explicitly addresses these gaps by proposing comprehensive, cognitively nuanced benchmarks directly aligned with the developmental data domain, thereby enabling accurate and relevant assessments of compact vision-language models trained from minimal, developmentally appropriate multimodal inputs.
@@ 3 Framework
Our proposed framework, BabyVLM, aims to facilitate resource-efficient pretraining and comprehensive evaluation of compact VLMs inspired by the minimal yet highly informative learning environments of human infants. To achieve this, BabyVLM comprises: (1) a filtered subset of baby-egocentric audio-visual recording from the SAYCam dataset, (2) a novel synthetic training dataset specifically crafted to reflect infant-directed linguistic and visual experiences, (3) a generative baseline model, BabyLLaVA, trained entirely on this developmentally plausible data, and (4) three novel evaluation benchmarks explicitly tailored to assess multimodal reasoning aligned with early cognitive stages, plus Labeled-S [31], an existing classification benchmark. Please refer to Figure 1 for an overview of our framework.
@@
These principles collectively ensure that the resulting BabyVLMs remain authentic representations of early-stage developmental models, with their effectiveness empirically confirmed in our evaluations (Section 4.5).
@@ 3.1 Datasets
Filtered SAYCam Dataset. The original SAYCam dataset [44] consists of egocentric audiovisual recordings. Following prior preprocessing by Vong et al. [50], we extract only child-directed speech and sample video into image-utterance pairs. To further align this dataset with developmental appropriateness, we refined the corpus by calculating CLIP similarity scores [35] between each image and its associated caption, retaining only pairs exceeding an experimentally determined similarity threshold of 0.2. This filtering ensures both linguistic and visual simplicity, consistent with the minimal complexity of early developmental stages, and results in approximately 67K image-utterance pairs. Examples of the filtered SAYCam dataset are provided in the supplementary material.
Figure 3: Illustrations of in-domain evaluation benchmarks in the BabyVLM framework. Labeled-S: The category label must be matched to the target referent among 4 candidates. Visual Two-Word Test: The positive phrase must be matched to the image. Positive and negative phrases are generated by GPT-4o. Baby Winoground: The positive and negative phrases must be matched with their corresponding images. Negative images are generated by Stable Diffusion, with prompts enhanced by GPT-4o. SAYCam Caption: The generated image caption must match the ground truth image caption. All image-caption pairs come from a de-duplicated subset of the SAYCam test split.
@@
Our approach includes two steps, as illustrated in Figure 2. In the first step, we utilize GPT-4o to rewrite original captions into simpler, child-directed utterances. Our prompting strategy explicitly instructed GPT-4o to produce concise, familiar, and straightforward sentences, emulating language that caregivers typically use with two-year-old children. At this step, we also prompt GPT-4o to identify which image-caption pairs are misaligned with an infant’s daily experience, and such examples are excluded from further processing. By emphasizing everyday vocabulary, simple grammar, and concrete objects and actions, we ensured linguistic alignment with early-stage learners.
In the second step, to further maintain visual consistency, we apply the Hungarian algorithm [17] and utilize CLIP similarity as a distance metric to select a subset of images resembing SAYCam. The number of selected samples matches the filtered SAYCam training set, resulting in a dataset that maintains visual alignment while balancing diversity and domain relevance. More details of the transformation guidelines and examples of rewritten captions are included in the supplementary material.
@@ 3.2 BabyLLaVA: Generative VLM Baseline
We then train a compact VLM, called BabyLLaVA, using the compiled dataset. Inspired by LLaVA [22], BabyLLaVA integrates a compact language model (GPT-2 [34], 7.18 million parameters) and vision encoder (ResNeXt-50 [54], 23 million parameters) through a lightweight multilayer perceptron connector. Consistent with our guiding principles, BabyLLaVA’s compact size and simplified architecture explicitly limit the model’s complexity, aligning closely with the realistic developmental constraints of early-stage learners. Training exclusively on our provided developmental dataset, BabyLLaVA provides a suitable baseline model to evaluate the effectiveness of developmentally aligned multimodal learning. Additional details are in the supplementary material.
@@ 3.3 Evaluation Tasks
To rigorously evaluate multimodal reasoning within the intended developmental scope, we introduce three novel benchmarks explicitly designed around cognitive milestones typical of early learners, in addition to the existing Labeled-S task. These tasks deliberately embody simplicity and developmental appropriateness, ensuring alignment with our guiding principles (Figure 3).
@@
Baby Winoground. Extending VTWT, Baby Winoground tests more advanced visio-linguistic compositional reasoning. Inspired by Winoground [48], this task presents two images and two corresponding phrases (one positive, one negative). Negative images were created by modifying specific visual elements of original images through targeted prompts provided to Stability AI’s Stable Image Ultra model [11]. All samples undergo a manual review to ensure minimal domain gaps and precise visio-linguistic mappings, resulting in 365 high-quality test samples. Full prompt-engineering details and validation methods appear in the supplementary material.
@@
SAYCam Caption. The SAYCam Caption benchmark evaluates generative captioning skills by requiring models to generate accurate, contextually relevant descriptions for SAYCam images. Captions are sourced from the test split of SAYCam, using child-directed utterances as ground truth. To refine the dataset, we deduplicate frames with identical utterances, retaining only those with the highest CLIP image-caption similarity. This yields 1598 distinct image-caption pairs, which are then manually verified, resulting in 294 final test samples. Evaluation is performed using the METEOR metric [5]. This task measures a model’s ability to generate coherent, semantically appropriate child-directed descriptions. Examples of test samples are provided in the supplementary material.
@@ 4 Experiments
Category Model Labeled-S Visual Two-Word Test Baby Winoground SAYCam Caption
@@
Upper Bound Models LLaVA-v1.5-7B 0.7400 0.7851 0.4274 0.6575 0.6301 0.1657
@@
Baby Models BabyLLaVA 0.4195 0.6252 0.0658 0.3890 0.2301 0.1379
@@
Table 3: Evaluation results of in-domain tasks, where a higher score indicates better performance. For Labeled-S, we use the same target and foil testing samples as [31] and report accuracy. For the Visual Two-Word Test, we report accuracy. For Baby Winoground, we report the group score for different context. For SAYCam Caption, we report the METEOR score.
Our experimental evaluation aims to compare VLM architectures and training paradigms within a developmentally plausible setting, investigate the effectiveness of our synthetic child-directed dataset, and perform a fine-grained analysis of compositional reasoning.
@@ 4.1 In-Domain Benchmark Results
Notably, CVCL—a contrastive model—consistently outperforms the generative BabyLLaVA model across most tasks. This observation aligns with existing literature [58, 47, 13, 12], suggesting that contrastive models may be better suited to discriminative tasks, possibly due to their direct objective of learning joint visual-textual alignment. However, generative models like BabyLLaVA demonstrate reasonable performance on simpler compositional tasks such as VTWT, indicating substantial potential for improvement on more sophisticated compositional tasks like Baby Winoground. In particular, Baby Winoground reveals a stark asymmetry: baby models perform above chance when reasoning from in-distribution (positive) context, but below chance from out-of-distribution (negative) context, highlighting a systematic failure under distribution shift. Moreover, generative captioning, measured by SAYCam Caption scores, remains challenging for all models, emphasizing the additional complexity inherent in generating full linguistic descriptions from minimal data.
Model Labeled-S Visual Two-Word Test Baby Winoground SAYCam Caption
@@
CVCL-ori 0.6086 0.6494 0.0932 0.5068 0.2246 N/A
CVCL-aug 0.5805 0.7021 0.2027 0.4657 0.4493 N/A
CVCL-aug-random 0.6023 0.6835 0.1068 0.4739 0.2958 N/A
BabyLLaVA-ori 0.4195 0.6252 0.0658 0.3890 0.2301 0.1379
BabyLLaVA-aug 0.5364 0.6933 0.0822 0.3726 0.3096 0.1592
BabyLLaVA-aug-random 0.5155 0.6553 0.0877 0.3616 0.2712 0.1778
Table 4: Ablation study of our transferred dataset on in-domain tasks. $X$-ori: Only SAYCam. $X$-aug: SAYCam + child-directed data. $X$-aug-random: SAYCam + random general purpose data.
@@ 4.2 Transferred Dataset Ablation
We next perform an ablation study (Table 4) comparing models trained under three scenarios: using our SAYCam data only (ori), SAYCam plus our synthetic, child-directed dataset (aug), and SAYCam augmented with randomly selected general-domain data of the same size as the transferred dataset (aug-random).
We observe clear performance improvements in CVCL and BabyLLaVA when using our carefully curated dataset compared to random augmentation, particularly on compositional reasoning tasks such as VTWT and Baby Winoground. These results indicate that explicitly adapting general-domain data to reflect the linguistic simplicity and visual content of infant environments significantly enhances the data efficiency and overall alignment of the resulting models. Notably, for Baby Winoground, training with the transferred dataset substantially improves performance in the negative context setting, despite a slight drop in the positive context score; while the randomly selected dataset also improves the negative context score, the gains are smaller, indicating that our transferred dataset is more effective in helping baby models generalize to broader domains. In contrast, improvements in generative captioning remain modest, suggesting that further refinements, such as enriching the linguistic variety or introducing narrative structures, could improve generative performance.

### v2

@@ Abstract
Human infants rapidly develop visual reasoning skills from minimal input, suggesting that developmentally inspired pretraining could significantly enhance the efficiency of vision-language models (VLMs). Although recent efforts have leveraged infant-inspired datasets like SAYCam, existing evaluation benchmarks remain misaligned—they are either too simplistic, narrowly scoped, or tailored for large-scale pretrained models. Additionally, training exclusively on SAYCam overlooks the broader, diverse input from which infants naturally learn. To address these limitations, we propose BabyVLM, a novel framework comprising diverse in-domain evaluation benchmarks and a synthetic training dataset created via child-directed transformations of existing datasets. We demonstrate that VLMs trained with our synthetic dataset achieve superior performance on BabyVLM tasks compared to models trained solely on SAYCam or general-purpose data of the SAYCam size. BabyVLM thus provides a robust, developmentally aligned evaluation tool and illustrates how compact models trained on carefully curated data can generalize effectively, opening pathways toward data-efficient vision-language learning paradigms.
@@ 1 Introduction
We propose a novel framework, BabyVLM, for data-efficient pretraining of vision-language models (VLMs). To this end we introduce methods for creating minimal yet naturalistic data—akin to the input human infants receive—as well as diverse in-domain evaluation benchmarks. By carefully curating the training data, we show that our method yields more robust, baby-like representations compared to training on general-purpose corpora, and can further serve as a template for resource-efficient model training in other specialized domains.
Challenges in Current VLM Training. Vision-Language Models have advanced rapidly in recent years [41, 52, 16, 62, 28], but these advancements often rely on massive datasets and prohibitively expensive computational resources. For instance, training large-scale models such as LLaMA [56] or LLaVA [28] can require thousands of GPU hours [15, 49]. Such demands pose fundamental barriers for independent researchers with limited resources, highlighting the need for more accessible pretraining methods.
@@
  • In-Domain Evaluation Tasks. We design three novel evaluation tasks derived from the SAYCam dataset. These tasks are tailored to reflect the cognitive and perceptual abilities typical of early human development, enabling diverse and meaningful evaluation of compact models trained on developmentally plausible data.
@@
Figure 2: Pipeline for generating the transferred dataset. Step 1: We prompt GPT-4o to check whether an input caption is describing something a child would see in daily life and transfer the original image captions into simpler, child-directed utterances. Step 2: We use the CLIP similarity score as a metric to represent the distance between two images, and then conduct Hungarian matching to select a small subset of the transferred dataset that is visually aligned with SAYCam images.
@@ 2 Related Work
Vision-Language Models. Large vision-language models (VLMs) [41, 28, 16, 52, 4, 31] have significantly advanced multimodal understanding by integrating visual and linguistic data for various tasks, including image captioning, visual question answering, and conversational interaction. Early influential models such as CLIP [41] leveraged contrastive learning paradigms, effectively aligning visual and textual representations within a unified embedding space. More recent generative frameworks, such as LLaVA series [28, 26, 27, 23, 67, 22, 24], have combined pretrained visual encoders [41, 64] with large language models [70, 3, 53, 17], enabling more advanced conversational interactions and multimodal generative capabilities. However, these models typically require extensive computational resources and large-scale datasets. In contrast, our approach specifically targets compact generative VLMs trained exclusively on developmentally plausible datasets, providing a framework to improve data efficiency and better align model training with cognitive development processes observed in human infants.
@@
Multimodal Benchmarks. Existing multimodal evaluation benchmarks can be broadly classified as general-purpose or developmentally inspired. General-purpose benchmarks, such as Visual Question Answering (VQA) [2, 65] and Winoground [55], evaluate advanced visio-linguistic integration but typically rely on large-scale, non-developmental datasets, rendering them unsuitable for compact models trained on limited developmental data. Conversely, developmentally inspired benchmarks—such as Labeled-S [37], ModelVsBaby [47], DevBench [51], and MEWL [18]—are more aligned with early cognitive processes but often limited to simplistic classification tasks or utilize data not fully reflective of the training domain. Our work explicitly addresses these gaps by proposing cognitively nuanced benchmarks directly aligned with the developmental data domain, thereby enabling accurate and relevant assessments of compact vision-language models trained from minimal, developmentally appropriate multimodal inputs.
@@ 3 Framework
Our proposed framework, BabyVLM, aims to facilitate resource-efficient pretraining and developmentally aligned evaluation of compact VLMs inspired by the minimal yet highly informative learning environments of human infants. To achieve this, BabyVLM comprises: (1) a filtered subset of baby-egocentric audio-visual recording from the SAYCam dataset, (2) a novel synthetic training dataset specifically crafted to reflect infant-directed linguistic and visual experiences, (3) a generative baseline model, BabyLLaVA, trained entirely on this developmentally plausible data, and (4) three novel evaluation benchmarks explicitly tailored to assess multimodal reasoning aligned with early cognitive stages, plus Labeled-S [37], an existing classification benchmark. Please refer to Figure 1 for an overview of our framework.
@@
These principles collectively ensure that the resulting BabyVLMs remain authentic representations of early-stage developmental models, with their effectiveness empirically confirmed in our out-of-domain task evaluations, which are presented in the supplementary material.
@@ 3.1 Datasets
Filtered SAYCam Dataset. The original SAYCam dataset [50] comprises egocentric audiovisual recordings. Following Vong et al. [57], we extract child-directed speech and sample video into image-utterance pairs. To improve data quality, we filter pairs using CLIP similarities [41], retaining only those above a 0.2 threshold to ensure high image-text relevance. This results in  67K pairs. Examples are provided in the supplementary material.
Figure 3: Illustrations of in-domain evaluation benchmarks in the BabyVLM framework. Labeled-S: The category label must be matched to the target referent among 4 candidates. Visual Two-Word Test: The positive phrase must be matched to the image. Positive and negative phrases are generated by GPT-4o. Baby Winoground: The positive and negative phrases must be matched with their corresponding images. Negative images are generated by Stable Diffusion [12], with prompts enhanced by GPT-4o. SAYCam Caption: The generated image caption must match the ground truth image caption. All image-caption pairs come from a de-duplicated subset of the SAYCam test split.
@@
Our approach comprises two steps, as illustrated in Figure 2. In the first step, we prompt GPT-4o to rewrite captions into concise, child-directed utterances, emulating speech typically used with two-year-olds. GPT-4o also flags which image-caption pairs are misaligned with an infant’s daily experience for exclusion. By emphasizing everyday vocabulary, simple grammar, and concrete objects and actions, we ensured linguistic alignment with early-stage learners.
In the second step, to further maintain visual consistency, we apply the Hungarian algorithm [19] and utilize the CLIP similarity as a distance metric to select a subset of images resembing SAYCam. The number of selected samples matches that of the filtered SAYCam training set, resulting in a dataset that maintains visual alignment while balancing diversity and domain relevance. More details of the transformation guidelines and examples of rewritten captions are included in the supplementary material.
@@ 3.2 BabyLLaVA: Generative VLM Baseline
We then train a compact VLM, called BabyLLaVA, using the compiled dataset. Inspired by LLaVA [28], BabyLLaVA integrates a compact language model (GPT-2 [40], 7M parameters) and vision encoder (ResNeXt-50 [61], 23M parameters) through a lightweight multilayer perceptron connector. To examine the impact of model capacity, we also provide a larger variant (Llama-1.1B [66] + ViT-L [11]). Consistent with our guiding principles, BabyLLaVA’s compact size and simplified architecture explicitly limit the model’s complexity, aligning closely with the realistic developmental constraints of early-stage learners. Training exclusively on our provided developmental dataset, BabyLLaVA provides a suitable baseline model to evaluate the effectiveness of developmentally aligned multimodal learning. Additional details can be found in the supplementary material.
@@ 3.3 Evaluation Tasks
To rigorously evaluate multimodal reasoning within the intended developmental scope, we introduce three novel benchmarks explicitly designed around cognitive milestones typical of early learners, in addition to the existing Labeled-S task [37]. These tasks deliberately embody simplicity and developmental appropriateness, ensuring alignment with our guiding principles (Figure 3).
@@
Baby Winoground. Extending VTWT, Baby Winoground tests more advanced visio-linguistic compositional reasoning. Inspired by Winoground [55], this task presents two images and two corresponding phrases (one positive, one negative). Negative images were created by modifying specific visual elements of original images through targeted prompts provided to Stability AI’s Stable Image Ultra model [12]. All samples undergo a manual review to ensure minimal domain gaps and precise visio-linguistic mappings, resulting in 365 high-quality test samples. Full prompt-engineering details and validation methods are in the supplementary material.
@@
SAYCam Caption. The SAYCam Caption benchmark evaluates generative captioning skills by requiring models to generate accurate, contextually relevant descriptions for SAYCam images. Captions are sourced from the test split of SAYCam, using child-directed utterances as ground truth. To refine the dataset, we deduplicate frames with identical utterances, retaining only those with the highest CLIP image-caption similarities. This yields 1598 distinct image-caption pairs, which are then manually verified, resulting in 294 final test samples. Evaluation is performed using the METEOR metric [5]. This task measures a model’s ability to generate coherent, semantically appropriate child-directed descriptions. Examples of test samples are provided in the supplementary material.
@@ 4 Experiments
Category Model Labeled-S VTWT Baby Winoground
SAYCam
Caption
@@
Upper Bound
Models
LLaVA-v1.5-7B 0.7400 0.7851 0.4274 0.6575 0.6301 0.1657
@@
Baby Models BabyLLaVA-GPT2 0.4195 0.6252 0.0658 0.3890 0.2301 0.1379
BabyLLaVA-Llama 0.4200 0.6029 0.0521 0.3945 0.2466 0.1287
@@
Table 3: Evaluation results of in-domain tasks, where a higher score indicates better performance. For Labeled-S, we use the same target and foil testing samples as [37] and report accuracy. For the Visual Two-Word Test (VTWT), we report accuracy. For Baby Winoground, we report the group score for different contexts. For SAYCam Caption, we report the METEOR score.
Our experimental evaluation aims to compare VLM architectures and training paradigms within a developmentally plausible setting, investigate the effectiveness of our synthetic child-directed dataset, and perform a fine-grained analysis of compositional reasoning. To validate that our baby models align with the cognitive and linguistic limitations of early-stage learners, we also assess baby models on tasks that exceed typical infant-level developmental capacities, presented in the supplementary materials.
@@ 4.1 In-Domain Benchmark Results
Notably, CVCL—a contrastive model—consistently outperforms the generative BabyLLaVA model across most tasks. This observation aligns with existing literature [68, 54, 14, 13], suggesting that contrastive models may be better suited to discriminative tasks, possibly due to their direct objective of learning joint visual-textual alignment. However, generative models like BabyLLaVA demonstrate reasonable performance on simpler compositional tasks such as VTWT, indicating substantial potential for improvement on more sophisticated compositional tasks like Baby Winoground. Interestingly, the larger BabyLLaVA-Llama variant performs similarly or even worse than BabyLLaVA-GPT2, despite being about 50 times larger—suggesting overfitting to the limited data; as such, we default to using the GPT2 version for all the remaining discussion. In particular, Baby Winoground reveals a stark asymmetry: baby models perform above chance when reasoning from in-distribution (positive) context, but below chance from out-of-distribution (negative) context, highlighting a systematic failure under distribution shift. Moreover, generative captioning, measured by SAYCam Caption scores, remains challenging for all models, emphasizing the additional complexity inherent in generating full linguistic descriptions from minimal data.
Model Labeled-S VTWT Baby Winoground
SAYCam
Caption
@@
CVCL-filtered 0.6086 0.6494 0.0932 0.5068 0.2246 N/A
CVCL-filtered-aug 0.5805 0.7021 0.2027 0.4657 0.4493 N/A
CVCL-filtered-random 0.6023 0.6835 0.1068 0.4739 0.2958 N/A
BabyLLaVA-filtered 0.4195 0.6252 0.0658 0.3890 0.2301 0.1379
BabyLLaVA-filtered-aug 0.5364 0.6933 0.0822 0.3726 0.3096 0.1592
BabyLLaVA-filtered-random 0.5155 0.6553 0.0877 0.3616 0.2712 0.1778
BabyLLaVA-filtered-double 0.4659 0.6383 0.0658 0.2411 0.2301 0.1799
BabyLLaVA-aug-only 0.5000 0.6239 0.0630 0.3370 0.3644 0.0615
BabyLLaVA-random-only 0.4400 0.5098 0.0548 0.2904 0.3836 0.0615
Table 4: Ablation study of our transferred dataset on in-domain tasks. $X$-filtered: Only SAYCam. $X$-filtered-aug: SAYCam + child-directed transferred data. $X$-filtered-random: SAYCam + random general-domain data. $X$-filtered-double: SAYCam with double-numbered samples. $X$-aug-only: Only random general-domain data. $X$-random-only: Only random general-domain data.
@@ 4.2 Transferred Dataset Ablation
We next perform an ablation study (Table 4) comparing models trained on several different dataset settings: using our filtered SAYCam dataset only (filtered-only), SAYCam plus our transferred, child-directed dataset (filtered-aug), transferred dataset only (aug-only), SAYCam plus randomly selected general-domain dataset of the same size as the transferred dataset (filtrered-random), random general-domain dataset only (random-only), and a bigger filtered SAYCam dataset whose number of samples is doubled by relaxing the threshold (filtered-double). Additional analysis regarding data efficiency can be found in the supplementary material.
We observe clear performance improvements in CVCL and BabyLLaVA when using our carefully curated dataset compared to random augmentation, particularly on compositional reasoning tasks such as VTWT and Baby Winoground. These results indicate that explicitly adapting general-domain data to reflect the linguistic simplicity and visual content of infant environments significantly enhances the data efficiency and overall alignment of the resulting models. Using only the transferred data degrades the model performance, which shows the importance of having the original SAYCam dataset as an anchor. Notably, for Baby Winoground, training with the transferred dataset substantially improves performance in the negative context setting, despite a slight drop in the positive context score; while the randomly selected dataset also improves the negative context score, the gains are smaller, indicating that our transferred dataset is more effective in helping baby models generalize to broader domains. In contrast, improvements in generative captioning remain modest, suggesting that further refinements, such as enriching the linguistic variety or introducing narrative structures, could improve generative performance.